# Euclidean Distance Matrix Completion via Asymmetric Projected Gradient Descent

### v1

@@ I-B Major Contributions
^$\hat{\nabla}f(\mathbf{P})=\frac{2}{p}\mathcal{R}_{\Omega}(\mathbf{P}\mathbf{P}^{T}-\mathbf{G}^{\star})\mathbf{P},$ (4b)
@@
which, can be viewed as optimizing $\mathbf{P}$ over a quotient manifold. Compared with [1], our major contributions are three fold:
@@
  • We analyze the contraction behavior when using (4b) to perform Projected GD (PGD), starting from a surrogate returned by a recent proposed spectral estimator[1][23, Lemma 5.6] called One-step MDS (OS-MDS). This routine is referred to as the Asymmetric Projected Gradient Descent (APGD), where the projection step enforces iteration to stay incoherent, similar to the setup in[35]. We show a near linear convergence toward the ground truth of the APGD iteration can be established given $\mathcal{O}(\mu^{2}r^{3}\kappa^{2}n\log n)$ random distance samples by constructing standard regularity condition w.r.t. the pseudo gradient direction.
@@
  • We provide two new uniform estimates for controlling two types of the non-tangent space residual during iteration in inner product form, namely, Lemma B.1 and B.2, thereby replacing the random graph lemma used in classic non-convex LRMC tasks. These bounds are developed based on a celebrated sharp result by Bandeira and van Handel[60], also the separation trick originated from Bhojanapalli and Jain[61], substantially different with the one used in our previous work[22, Lemma B.2].
@@
  • We present numerical comparison between the OS-MDS initialized un-regularized s-stress (1) and APGD with vanilla Barzilai-Borwein stepsize over synthetic point set. While the APGD is poorly behaved and appears to possess practical significance only in the near-asymptotic region, phenomena analogous to the role of incoherence restrictions in non-convex LRMC[62, Ch. 2] are observed. This raises interesting questions on the design principle of non-convex matrix recovery strategies under non-orthonormal, non-sub-Gaussian, and ill-conditioned sample basis – is EDMC a special case?
@@
Organization: In Section II we provide a more detailed background and model setup. We state our main theorem and two important lemmas that control the initialization error of OS-MDS and contraction behavior of APGD in Section II-A. The proofs are referred to the Appendix. Section III contains the aforementioned numerical result and the manuscript is concluded with a discussion in Section IV.
@@
Notation: $\mathbf{y}$, $\mathbf{Y}$ stands for column vectors and matrices. Superscript $(\cdot)^{\star}$ emphasize the ground truth. $\mathbf{1}$, $\mathbf{0}$, $\mathbf{I}$ is the all ones, all zeros vector, and the identity matrix. $\mathbf{Y}_{ij}$, $\mathrm{diag}(\mathbf{Y})$, and $\mathrm{diag}(\mathbf{y})$ is the $(i,j)$-th entity, column vector formed by the diagonal elements of $\mathbf{Y}$, and diagonal matrix formed by $\mathbf{y}$, respectively. $\mathbf{Y}_{\cdot,k}$ / $\mathbf{Y}_{k,\cdot}^{T}$ stands for the $k$-th column/row of $\mathbf{Y}$. $\mathrm{Od}(\mathbf{Y})$ means a matrix obtained by nullifying the diagonal of $\mathbf{Y}$. $\mathcal{F}^{*}$ stands for the adjoint of a linear operator $\mathcal{F}$, and $\mathcal{I}$ denotes the identity operator. $\circ$ is the Hadamard product. $\|\mathbf{Y}\|$, $\|\mathbf{Y}\|_{F}$, $\|\mathbf{Y}\|_{\infty}$, $\|\mathbf{Y}\|_{2,\infty}$, $\|\mathbf{Y}\|_{*}$ stand for the spectral, Frobenius, entry-wise $l_{\infty}$, row-wise $l_{2,\infty}$, and nuclear norm of $\mathbf{Y}$. $\mathcal{S}(n)$ denotes the set of symmetric matrix in $\mathbb{R}^{n\times n}$, and $O(r)$ stands for the $r$ dimensional orthogonal group. $\|\mathbf{y}\|_{2}$, $\|\mathbf{y}\|_{\infty}$ is the vector $l_{2}$, $l_{\infty}$ norm. $C,\,c>0$ are universal constants that may differ from line to line. We retain $\beta>2$ for expression like $n^{1-\beta}$, and $C_{\beta}$ denotes a constant that dependent with $\beta$. $a\lesssim b$ and $a\gtrsim b$ stands for $a<Cb$ and $a>Cb$, respectively. $\mathbf{Y}\succcurlyeq\mathbf{0}$ means that $\mathbf{Y}$ is PSD.
@@ II Mathematical Setup and Main Result
Let $\mathbf{P}^{\star}=[\mathbf{p}_{1}^{\star},\dots,\mathbf{p}_{n}^{\star}]^{T}\in\mathbb{R}^{n\times r}$ be the ground truth point set relative position and assume it has full column rank. The following non-convex QCQP (Quadratically Constrained Quadratic Programming) is usually referred to be the EDMC problem
$\mathrm{find}$ $\mathbf{G}$ (5)
$\mathrm{s.t.}$ $\mathbf{G}_{ii}+\mathbf{G}_{jj}-2\mathbf{G}_{ij}=\mathbf{D}^{\star}_{ij},\,(i,j)\in\Omega,$
$\mathbf{G1}=\mathbf{0},\,\mathbf{G}\succcurlyeq\mathbf{0},\,\mathrm{rank}(\mathbf{G})=r.$
Where $\mathbf{D}^{\star}_{ij}=d_{ij}^{\star 2}=\|\mathbf{p}_{i}^{\star}-\mathbf{p}_{j}^{\star}\|_{2}^{2}=\langle\mathbf{P}^{\star}\mathbf{P}^{\star T},\boldsymbol{\omega}_{\boldsymbol{\alpha}}\rangle$ denotes the $(i,j)$-th element in the ground truth EDM, and $\boldsymbol{\omega}_{\boldsymbol{\alpha}}=(\mathbf{e}_{i}-\mathbf{e}_{j})(\mathbf{e}_{i}-\mathbf{e}_{j})^{T}$ is the so-called EDM basis. $\Omega\subset\mathbb{I}:=\{(i,j):i\leq i<j\leq n\}$ [^6] Since the EDM is a symmetric and hollow matrix, we can omit its diagonal and only sample the upper triangle sub-matrix, which is referred to as the symmetric sampling later., and $\mathbf{G}$ denotes the Gram matrix formed by $\mathbf{P}\mathbf{P}^{T}$. Obviously, the absolute configuration of points is lost and we can only recover $\mathbf{P}^{\star}$ up to translation, rotation, and reflection. The translation is removed by only considering self-centered point set [^7] W.l.o.g., we can always assume $\mathbf{P}^{\star T}\mathbf{1}=\mathbf{0}$, and all considered point sets are in $\mathcal{S}_{c}^{n}$. That is, the identity operator $\mathcal{I}$ used in this work actually denotes $\mathcal{I}_{\{\mathbf{1}\}^{\perp}}$. We abuse =$\mathcal{I}=\mathcal{I}_{\{\mathbf{1}\}^{\perp}}$ for notation simplicity, one may also use different geometry centers, c.f.,[63].. The rotation and reflection ambiguity gives rise to a group synchronization problem over $O(r)$, and we use the solution of the orthogonal Procrustes problem to define the distance between $\mathbf{P}$ and $\mathbf{P}^{\star}$
@@
and the solution of (6) is denoted as $\boldsymbol{\psi}^{\star}$. It satisfies $\mathbf{P}^{T}\mathbf{P}^{\star}\boldsymbol{\psi}^{\star}\succcurlyeq\boldsymbol{0}$ and $\Delta^{T}\mathbf{P}^{\star}\boldsymbol{\psi}^{\star}\in\mathcal{S}(r)$[54, Lemma 6]. We also let $\Delta$ and $\boldsymbol{\psi}^{\star}$ inherit any super/sub-script of $\mathbf{P}$, i.e., $\mathrm{dist}(\mathbf{P}^{(\cdot)},\mathbf{P}^{\star})=\|\Delta^{(\cdot)}\|_{F}=\|\mathbf{P}^{(\cdot)}-\mathbf{P}^{\star}\boldsymbol{\psi}^{\star(\cdot)}\|_{F}$.
@@
A list of works[42][63][64] shows that by restricting $\mathcal{I}$ onto the centered subspace $\mathcal{S}_{c}^{n}:=\{\mathbf{A}:\mathbf{A1}=\mathbf{0},\mathbf{A}\in\mathcal{S}(n)\}$ and form $\mathcal{I}_{\{\mathbf{1}\}^{\perp}}$, the following decomposition holds over $\mathcal{S}_{c}^{n}$
@@
and $\mathcal{I}_{\{\mathbf{1}\}^{\perp}}(\mathbf{A})=\mathbf{A}$ if $\mathbf{A}\in\mathcal{S}_{c}^{n}$. They call $\boldsymbol{\nu}_{\boldsymbol{\alpha}}:=-\frac{1}{2}\mathbf{J}(\mathbf{e}_{i}\mathbf{e}_{j}^{T}+\mathbf{e}_{j}\mathbf{e}_{i}^{T})\mathbf{J}$ the dual-basis of $\boldsymbol{\omega}_{\boldsymbol{\alpha}}$[42]. From (10), we find the explicit expression of the pre-conditioned sample operator $\mathcal{R}_{\Omega}$
@@
Under the symmetric Bernoulli model with parameter $p$, we have $\mathbb{E}\frac{1}{p}\mathcal{R}_{\Omega}=\mathbb{E}\frac{1}{p}\mathcal{R}_{\Omega}^{*}=\mathcal{I}$. Finally, let $\mathcal{Q}_{\Omega}$ be defined as
$\mathcal{Q}_{\Omega}(\cdot):=\mathcal{P}_{\Omega}g(\cdot)=\sum_{\boldsymbol{\alpha}\in\mathbb{I}}\delta_{\boldsymbol{\alpha}}\langle\cdot,\boldsymbol{\omega}_{\boldsymbol{\alpha}}\rangle(\mathbf{e}_{i}\mathbf{e}_{j}^{T}+\mathbf{e}_{j}\mathbf{e}_{i}^{T}).$ (12)
@@ II-A Main Result
Algorithm 1 APGD
0:  Pseudo-gradient ^$\hat{\nabla}f(\mathbf{P})$ as in (4b), sampled distances $\mathcal{P}_{\Omega}\mathbf{D}^{\star}$.
1:  Initial APGD by OS-MDS as in [1][23]
$\hat{\mathbf{P}}\hat{\mathbf{P}}^{T}=\mathcal{T}_{r}\left[-\frac{1}{2p}\mathbf{J}(\mathcal{P}_{\Omega}\mathbf{D}^{\star})\mathbf{J}\right],\,\mathbf{P}_{1}=\mathcal{P}_{I}(\hat{\mathbf{P}}),$
where $\mathcal{P}_{I}$ is defined by (14).
2:  for $k=1,2,\dots$ do
4:     $\mathbf{P}_{k+1}=\mathcal{P}_{I}(\mathbf{P}_{k+1})$.
5:  end for
6:  return  $\mathbf{P}_{k}$
@@
holds with probability at least $1-cn^{1-\beta}-n^{2-\beta}$. Where $0<\eta^{\prime}\leq\frac{1}{336c_{I}\mu r}$, $c_{c}=\frac{80}{7}$, and $\Delta_{k}$ refers substituting $\mathbf{P}_{k}$ into (6). To reach $\varsigma$ stationary point, i.e., $\|\Delta_{k}\|_{F}^{2}/\|\Delta_{0}\|_{F}^{2}\leq\varsigma$, APGD needs around $336c_{I}\mu r\log(\frac{1}{\varsigma})$ iterations.
@@
:=$\mathcal{B}:=\{\mathbf{P}\,:\,\|\Delta\|_{2,\infty}\leq\sqrt{\frac{c_{I}\mu r\sigma_{1}^{\star}}{n}},\|\Delta\|_{F}^{2}\leq\frac{\sigma_{r}^{\star}}{c_{c}}\},\,c_{I}>4.$
@@
$\hat{\mathbf{P}}\hat{\mathbf{P}}^{T}=\mathcal{T}_{r}\left[-\frac{1}{2p}\mathbf{J}(\mathcal{P}_{\Omega}\mathbf{D}^{\star})\mathbf{J}\right],\,\mathbf{P}_{1}=\mathcal{P}_{I}(\hat{\mathbf{P}}),$
@@
satisfies
$\|\Delta_{k+1}\|_{F}^{2}\leq(1-\frac{\eta\sigma_{r}^{\star}}{2})\|\Delta_{k}\|_{F}^{2},$
with probability at least $1-cn^{1-\beta}-n^{2-\beta}$, provided with $p\gtrsim C_{\beta}\frac{\mu^{2}r^{2}\kappa^{2}\log n}{n}$, and $0<\eta\leq\frac{1}{168\sigma_{1}^{\star}c_{I}\mu r\kappa}$.

### v2

@@ I-C Major Contributions and Organization
$\hat{\nabla}f(\mathbf{P})=\frac{2}{p}\mathcal{R}_{\Omega}(\mathbf{P}\mathbf{P}^{T}-\mathbf{G}^{\star})\mathbf{P},$ (7b)
@@
which, can be viewed as optimizing $\mathbf{P}$ over a quotient manifold. Compared with [1], our major contributions are twofold:
@@
  • We analyze the contraction behavior when using (7b) to perform Projected Gradient Descent (PGD), starting from a surrogate returned by a recently proposed spectral estimator[1][23, Lemma 5.6], called One-step MDS (OS-MDS). This routine is referred to as the Asymmetric Projected Gradient Descent (APGD), where the “projection” step enforces iteration to stay incoherent, similar to the setup in[38]. And the name “asymmetric” comes from the fact that linear sensing operator $\mathcal{R}_{\Omega}$ is not self-adjoint, necessitating a delicate analysis of different components of the iteration residual, a feature not typically encountered when dealing with restricted strong convexity w.r.t. the first-order derivative. We show a near-linear convergence to the ground truth of the APGD iteration can be established given $\mathcal{O}(\mu^{2}r^{3}\kappa^{2}n\log n)$ random distance samples by constructing standard regularity condition w.r.t. the pseudo gradient direction.
@@
  • We provide two new uniform estimates for controlling two types of the non-tangent space residuals during iteration in inner product form, namely, Lemma B.1 and B.2, which act as analogies and extensions of the random graph lemma used in classic non-convex LRMC tasks[36, Lemma 7.1][41, Lemma 5]. These bounds are developed based on a celebrated sharp result by Bandeira and van Handel[58], as well as the separation trick originated from Bhojanapalli and Jain[59] in deterministic matrix completion context with Ramanujan graphs, substantially different from the one used in our previous work[22, Lemma B.2]. Moreover, these two estimates match the best known bound in incoherent LRMC for resemblance purposes[60, Lemma 8] up to $\log n$ factor [^9] Readers interested in comparing these non-tangent space concentration results are referred to [60, Sec. 4.1].. Rationales behind our choice of analyzing local behavior of (7b), and the current dilemma between analytical tractability and algorithmic fidelity behind this specific attempt, are explained in Section II-C.
@@
We also draw numerical comparisons between the OS-MDS initialized unregularized s-stress (3), IFHT-EDMC[23, Alg. 2], and APGD with vanilla Barzilai-Borwein (BB) stepsize over synthetic point set. While the APGD is poorly behaved and appears to possess practical significance only in the near-asymptotic region, phenomena analogous to the role of incoherence restrictions in non-convex LRMC[61, Ch. 2] are observed. This raises interesting questions regarding the design principles of non-convex matrix recovery strategies under non-orthonormal, non-sub-Gaussian, and ill-conditioned sample basis – is EDMC a special case?
This section concludes with a comparison of the theoretical results presented in this work with those from three recent studies[20][23][31], as summarized by Table I. While similarities lie in the assumption of incoherence restrictions, uniform sampling, and the use of dual basis expansion, their algorithmic frameworks vary. The term “strong incoherence” is referred to either[32, Eq. (3)] or[20, Def. 1], while the former being known to be unnecessary in LRMC [^10] We omit extra $\mathcal{O}(\mathrm{poly}(r))$ factor that may implicitly contained by the strong incoherence parameter when drawing the third column of Table I., reducing the latter in modified Golfing scheme for EDMC remains an open question.
@@
Organization: In Section II we provide a more detailed background and model setup. We state our main theorem and two important lemmas that control the initialization error of OS-MDS and contraction behavior of APGD in Section II-A. The proofs are deferred to the Appendix. Section III contains the aforementioned numerical result and the manuscript is concluded with a discussion in Section IV.
@@
Notation: $\mathbf{y}$, $\mathbf{Y}$ stands for column vectors and matrices. $\mathbf{Y}_{ij}$, $\mathrm{diag}(\mathbf{Y})$, and $\mathrm{diag}(\mathbf{y})$ is the $(i,j)$-th entity, column vector formed by the diagonal elements of $\mathbf{Y}$, and diagonal matrix formed by $\mathbf{y}$, respectively. $\mathbf{Y}_{\cdot,k}$ / $\mathbf{Y}_{k,\cdot}^{T}$ stands for the $k$-th column/row of $\mathbf{Y}$. $\mathrm{Od}(\mathbf{Y})$ means a matrix obtained by nullifying the diagonal of $\mathbf{Y}$. $\mathcal{F}^{*}$ stands for the adjoint of a linear operator $\mathcal{F}$, and $\mathcal{I}$ denotes the identity operator. $\circ$ is the Hadamard product. $\|\mathbf{Y}\|$, $\|\mathbf{Y}\|_{F}$, $\|\mathbf{Y}\|_{\infty}$, $\|\mathbf{Y}\|_{2,\infty}$, $\|\mathbf{Y}\|_{*}$ stand for the spectral, Frobenius, entry-wise $l_{\infty}$, row-wise $l_{2,\infty}$, and nuclear norm of $\mathbf{Y}$. $O(r)$ stands for the $r$ dimensional orthogonal group. $\|\mathbf{y}\|_{2}$, $\|\mathbf{y}\|_{\infty}$ is the vector $_{2}, ll_{\infty}$ norm. C,c>0C,\,c>0 are universal constants that may differ from line to line. We retain $\beta>2$ for expression like $n^{1-\beta}$, and $C_{\beta}$ denotes a constant that dependent with $\beta$ and some $c_{I}>4$. $\epsilon$ and $\varepsilon\leq 1$ are used to denote absolute numbers that control finite sample error. $a\lesssim b$ and $a\gtrsim b$ stands for $a<Cb$ and $a>Cb$, respectively. $\mathbf{Y}\succcurlyeq\mathbf{0}$ means that $\mathbf{Y}$ is PSD. $\mathbb{S}^{n-1}$ stands for the unit sphere in $\mathbb{R}^{n}$ centered at origin.
@@ II Mathematical Setup and Main Result
Without loss of generality, we can always assume $\mathbf{P}^{\star T}\mathbf{1}=\mathbf{0}$, and all considered point sets are in the centered subspace $\mathcal{S}_{c}^{n}:=\{\mathbf{A}:\mathbf{A1}=\mathbf{0},\mathbf{A}\in\mathcal{S}(n)\}$ to remove the translation ambiguity. For notation simplicity, we abuse $\mathcal{I}$ to denote its $\mathcal{S}_{c}^{n}$ restriction $\mathcal{I}_{\{\mathbf{1}\}^{\perp}}$, i.e., $\mathcal{I}_{\{\mathbf{1}\}^{\perp}}(\mathbf{A})=\mathbf{A}$ if $\mathbf{A}\in\mathcal{S}_{c}^{n}$. One may also use different geometry centers[62]. The rotation and reflection ambiguity in EDMC gives rise to a group synchronization problem over $O(r)$, and we use the solution of the orthogonal Procrustes problem to define the distance between $\mathbf{P}$ and $\mathbf{P}^{\star}$
@@
and the solution of (8) is denoted as $\boldsymbol{\psi}^{\star}$. It satisfies $\mathbf{P}^{T}\mathbf{P}^{\star}\boldsymbol{\psi}^{\star}\succcurlyeq\boldsymbol{0}$ and $\Delta^{T}\mathbf{P}^{\star}\boldsymbol{\psi}^{\star}\in\mathcal{S}(r)$[63, Lemma 6]. We also let $\Delta$ and $\boldsymbol{\psi}^{\star}$ inherit any super/sub-script of $\mathbf{P}$, i.e., $\mathrm{dist}(\mathbf{P}^{(\cdot)},\mathbf{P}^{\star})=\|\Delta^{(\cdot)}\|_{F}=\|\mathbf{P}^{(\cdot)}-\mathbf{P}^{\star}\boldsymbol{\psi}^{\star(\cdot)}\|_{F}$.
@@
A list of works[48][62][64] shows that the following primal-dual basis decomposition of $\mathcal{I}_{\{\mathbf{1}\}^{\perp}}$ holds over $\mathcal{S}_{c}^{n}$
@@
and they call $\boldsymbol{\nu}_{\boldsymbol{\alpha}}:=-\frac{1}{2}\mathbf{J}(\mathbf{e}_{i}\mathbf{e}_{j}^{T}+\mathbf{e}_{j}\mathbf{e}_{i}^{T})\mathbf{J}$ the dual-basis of $\boldsymbol{\omega}_{\boldsymbol{\alpha}}$[48]. From (12), we find the explicit expression of the pre-conditioned sample operator $\mathcal{R}_{\Omega}$
@@
Under the symmetric Bernoulli model with parameter $p$, we have $\mathbb{E}\frac{1}{p}\mathcal{R}_{\Omega}=\mathbb{E}\frac{1}{p}\mathcal{R}_{\Omega}^{*}=\mathcal{I}$. Finally, let $\mathcal{Q}_{\Omega}$ be the sensing operator in (3), that is
$\mathcal{Q}_{\Omega}(\cdot):=\mathcal{P}_{\Omega}g(\cdot)=\sum_{\boldsymbol{\alpha}\in\mathbb{I}}\delta_{\boldsymbol{\alpha}}\langle\cdot,\boldsymbol{\omega}_{\boldsymbol{\alpha}}\rangle(\mathbf{e}_{i}\mathbf{e}_{j}^{T}+\mathbf{e}_{j}\mathbf{e}_{i}^{T}),$ (14)
and we abuse
$\tilde{\mathcal{Q}}_{\Omega}(\cdot):=\sum_{\boldsymbol{\alpha}\in\mathbb{I}}\delta_{\boldsymbol{\alpha}}\langle\cdot,\boldsymbol{\omega}_{\boldsymbol{\alpha}}\rangle\mathbf{e}_{\boldsymbol{\alpha}},\,\tilde{\mathcal{Q}}_{\Omega}:\mathbb{R}^{n\times n}\to\mathbb{R}^{L},$
to denote the re-scaled vectorization of (14), where $\mathbf{e}_{\boldsymbol{\alpha}}$ is the $\boldsymbol{\alpha}$-th canonical Euclidean basis in $\mathbb{R}^{L}$. It is straightforward to show $\mathcal{Q}_{\Omega}^{*}\mathcal{Q}_{\Omega}=2\tilde{\mathcal{Q}}_{\Omega}^{*}\tilde{\mathcal{Q}}_{\Omega}$.
@@ II-A Main Result
holds with probability at least $1-cn^{1-\beta}-n^{2-\beta}$. Where $0<\eta^{\prime}\leq\frac{1}{336c_{I}\mu r\kappa^{2}}$, $c_{c}=\frac{80}{7}$, and $\Delta_{k}$ refers substituting $\mathbf{P}_{k}$ into (8). To reach $\varsigma$ stationary point, i.e., $\|\Delta_{k}\|_{F}^{2}/\|\Delta_{0}\|_{F}^{2}\leq\varsigma$, APGD needs around $336c_{I}\mu r\kappa^{2}\log(\frac{1}{\varsigma})$ iterations.
@@
$\mathcal{B}:=\{\mathbf{P}\,:\,\|\Delta\|_{2,\infty}\leq\sqrt{\frac{c_{I}\mu r\sigma_{1}^{\star}}{n}},\|\Delta\|_{F}^{2}\leq\frac{\sigma_{r}^{\star}}{c_{c}}\},\,c_{I}>4.$
@@
[…]
@@
Algorithm 1 APGD
0: Pseudo-gradient $\hat{\nabla}f(\mathbf{P})$ as in (7b), sampled distances $\mathcal{P}_{\Omega}\mathbf{D}^{\star}$.
1: Initial APGD by OS-MDS as in [1][23]
[…]
where $\mathcal{P}_{I}$ is defined by (16).
2:for $k=1,2,\dots$ do
4:  $\mathbf{P}_{k+1}=\mathcal{P}_{I}(\mathbf{P}_{k+1})$.
5:end for
6:return $\mathbf{P}_{k}$
@@
satisfies $\|\Delta_{k+1}\|_{F}^{2}\leq(1-\frac{\eta\sigma_{r}^{\star}}{2})\|\Delta_{k}\|_{F}^{2}$ with probability at least $1-cn^{1-\beta}-n^{2-\beta}$, provided with $p\gtrsim C_{\beta}\frac{\mu^{2}r^{2}\kappa^{2}\log n}{n}$, and $0<\eta\leq\frac{1}{168\sigma_{1}^{\star}c_{I}\mu r\kappa}$.